PRELIMINARY ANALYSIS OF A C-TYPE BAND IN THE HIGH-RESOLUTION INFRARED SPECTRUM OF EE-1,4-DIFLUOROBUTADIENE

\underline{NORMAN C. CRAIG} and LAURA PEDRAZA, Department of Chemistry, Oberlin College, Oberlin, OH 44074; MICHAEL LOCK, Physikalisch-Chemisches-Institut der Justus-Liebig-Universität, D-35392 Giessen, Germany.

As part of a study of nonpolar molecules that exhibit surprising energy relationships among isomers, a C-type band in the infrared spectrum of the EE (trans-trans) isomer of 1,4-difluorobutadiene is being investigated. This band, centered at 934 cm^{−1}, is due to CH out-of-plane flapping. Although the band structure contains evidence of substantial perturbations, the rotational structure at 0.002 cm^{−1} resolution appears to be fully assignable. Preliminary rotational constants for the ground state are A = 1.0507499 (14), B = 0.0389603 (47), and C = 0.0375881 (40) cm^{−1} for this near-prolate symmetric top with κ = -0.99729. A synthetic method has been found to make the deuterated isotopomers needed for the complete structural analysis of the isomers of 1,4-difluorobutadiene.